OpenStudy (anonymous):

integral tan^5 4x dx

zepdrix (zepdrix):

Hey Biz :) Hmmmm an odd power of tangent. Thinkinggg...... \[\Large\rm \int\limits \tan^54x~dx=\int\limits (\tan4x)^2 \tan^34x~dx\]\[\Large\rm \]\[\Large\rm \int\limits (\sec^24x-1)\tan^34x~dx\]\[\Large\rm =\int\limits \tan^34x~\sec^24x~dx-\int\limits \tan^34x~dx\]In the first integral you can make the substitution: \(\Large\rm u=tan4x\) And it should work out nicely since your du will involve a sec^2(4x).

zepdrix (zepdrix):

For the second integral, you'll have to change some more tangents into secants D:

zepdrix (zepdrix):

\[\Large\rm \int\limits \tan^34x~dx=\int\limits (\sec^24x-1)\tan 4x~dx\]\[\Large\rm =\int\limits \tan4x~\sec^24x~dx-\int\limits \tan4x~dx\]

zepdrix (zepdrix):

We get another integral that's easy to u-sub. And then you'll want to recall your tangent integral. If you don't remember it, you can write it as sin4x/cos4x and do u=cos4x.

OpenStudy (anonymous):

Thnx :)